# Attractive Metadata Attack: Inducing LLM Agents to Invoke Malicious Tools

### v1

@@ Abstract
Large language model (LLM) agents have demonstrated remarkable capabilities in complex reasoning and decision-making by leveraging external tools. However, this tool-centric paradigm introduces a previously underexplored attack surface: adversaries can manipulate tool metadata—such as names, descriptions, and parameter schemas—to influence agent behavior. We identify this as a new and stealthy threat surface that allows malicious tools to be preferentially selected by LLM agents, without requiring prompt injection or access to model internals. To demonstrate and exploit this vulnerability, we propose the Attractive Metadata Attack (AMA), a black-box in-context learning framework that generates highly attractive but syntactically and semantically valid tool metadata through iterative optimization. Our attack integrates seamlessly into standard tool ecosystems and requires no modification to the agent’s execution framework. Extensive experiments across ten realistic, simulated tool-use scenarios and a range of popular LLM agents demonstrate consistently high attack success rates (81%–95%) and significant privacy leakage, with negligible impact on primary task execution. Moreover, the attack remains effective even under prompt-level defenses and structured tool-selection protocols such as the Model Context Protocol, revealing systemic vulnerabilities in current agent architectures. These findings reveal that metadata manipulation constitutes a potent and stealthy attack surface, highlighting the need for execution-level security mechanisms that go beyond prompt-level defenses.
@@ 1 Introduction
Recent progress has shown that large language model (LLM) agents excel at executing diverse tasks, particularly when equipped with external tools for complex decision making and environmental interaction. Representative systems such as ReAct [24], ToolBench [18], and ToolCoder [5] unify planning, flexible tool-calling, and deep integration of heterogeneous resources, thereby markedly improving the automation and generalization abilities of LLM agents in domains including financial analysis [26], healthcare [1], and e-commerce [32, 22].
However, the tight coupling between LLM agents and an open tool ecosystem is giving rise to a new class of behavioral security risks [30]. Beyond traditional concerns over content safety [17, 21]—e.g., preventing the model from outputting sensitive or harmful information—adversaries can tamper with the tool-calling chain by manipulating tool outputs [31, 27, 12], crafting abnormal execution paths [20], or injecting malicious prompts [6, 28]. Such interventions indirectly steer an agent’s decision process, leading to information leakage or task mis-execution. For instance, a prompt-injected agent may inadvertently call a tool and reveal private data, or may corrupt user inputs in an e-commerce or medical scenario, causing severe consequences.
These attacks primarily rely on the direct alteration of prompts or the tool-calling chain [25]. In contrast, we identify a subtler yet powerful attack surface: manipulating the metadata (e.g., name, description, and parameter schema) of a malicious tool to induce an LLM agent to invoke it. We formalize this new threat as the Attractive Metadata Attack (AMA) (see Figure 1). AMA exploits the fact that current LLM agents typically determine tool invocation based on the user’s query, the task context, and the metadata associated with each available tool. In this context, an attacker can legitimately craft the tool’s metadata, imbuing it with properties that make it disproportionately attractive to the agent, thereby increasing its likelihood of being preferred over benign tools. This attack requires neither modifying the prompt template nor accessing model internals, yet it enables long-term and stealthy control over the agent’s behavior.
@@
To achieve the attack objective of AMA, we formulate the generation of high-inducement tool metadata as a state–action–value optimization problem, leveraging in-context learning from LLM [13, 4]. This optimization process is supported by constrained mechanisms that systematically mine and efficiently construct the tool metadata required for AMA, including generation traceability, weighted value evaluation, and batch generation. The generation pipeline iteratively produces tool metadata that significantly increases the selection probability of a malicious tool in both domain-specific and general scenarios. During an actual attack, the adversary simply replaces a benign tool’s original metadata with AMA-optimized metadata, thereby granting it the ability to deceptively entice agent invocation. Since the metadata remains syntactically and semantically valid, this procedure does not disrupt the agent’s execution pipeline at a structural level, thereby forming a stealthy, widely applicable, and detection-resistant threat.
Extensive experiments across ten canonical tool-calling scenarios and four mainstream LLM agents—including both open-source and commercial models—demonstrate the effectiveness of AMA, with attack success rates consistently ranging from 81% to 95%. The crafted malicious tools are frequently and covertly invoked by agents, resulting in significant privacy leakage while leaving the primary task flow virtually unaffected. Notably, the attack remains effective even under prompt-level defenses and structured tool-selection protocols such as the Model Context Protocol (MCP) [3], revealing systemic vulnerabilities in current agent workflows. Moreover, existing prompt-level defenses [29] are largely ineffective against metadata-based manipulation, underscoring the stealth, generalizability, and robustness of the proposed method.
@@
  • We formulate metadata crafting as a state–action–value optimization problem, using LLMs’ in-context learning to generate metadata that maximizes malicious tool invocation likelihood. And to support efficient and scalable metadata generation, we introduce three key mechanisms: generation traceability, weighted value evaluation, and batch generation.
  • We demonstrate the effectiveness of AMA across ten tool-use scenarios and four representative LLM agents. AMA achieves 81%–95% attack success rates with minimal task disruption and notable privacy leakage. It bypasses prompt-level defenses and structured protocols such as MCP, revealing systemic vulnerabilities in current agent architectures.
@@ 2 Related Work
While tool augmentation significantly enhances the action capabilities of LLM agents, it also introduces new and diverse attack surfaces [25, 30]. Recent studies demonstrate that adversaries can inject carefully crafted prompts [28] or subtly manipulated instructions to mislead the agent into invoking malicious tools, resulting in privacy leakage, behavioral manipulation, or resource misuse [6, 7]. Such attacks often rely on modifying the prompt template, or intermediate instructions, and are therefore typically detectable by prompt-level sanitization or instruction filtering [29].
Another line of work explores tool-side threats, showing that tampering with third-party API outputs can misguide agent behavior or cause unintended actions [31, 27]. More advanced attacks dynamically construct malicious command sequences within the tool chain, leveraging benign tool outputs to craft downstream payloads targeting attacker-controlled services [12]. Multi-stage adversarial pipelines further combine tool injection and input manipulation to capture queries, redirect data, or disrupt planning, leading to privacy breaches or unauthorized tool use [20]. Overall, these approaches exploit crafted tool outputs as the primary vector for triggering malicious behaviors.
Unlike prior approaches that rely on prompt injection, contextual tampering, or toolchain manipulation, our work focuses on constructing malicious tools that can be autonomously selected by the LLM agent under normal instructions. This creates a “silent” attack vector, wherein the agent willingly invokes attacker-supplied tools based solely on their crafted metadata. Such attacks require no prompt interference, yet achieve persistent influence by exploiting the agent’s internal tool-selection mechanism. As a result, AMA bypasses prompt-level defenses [29] entirely and reveals deeper structural vulnerabilities in current agent-tool interaction protocols [10].
@@ 3.1 Preliminaries and Settings
We consider an LLM-based agent that augments its reasoning and decision-making for complex tasks by invoking external tools. Let $\mathcal{T}$ denote a fixed toolset, where each tool $t\in\mathcal{T}$ is characterized by its name, description, and parameter schema. Upon receiving a user query (or task) $q$, we approximate the agent’s preference over tools via a latent scoring function: $t^{*}=\arg\max_{t\in\mathcal{T}}\mathcal{S}(q,\mathcal{O},P_{\mathrm{sys}},\mathrm{Meta}(t))$, where $\mathcal{S}(\cdot)$ denotes an implicit score function that captures the agent’s inclination to invoke tool $t$ under the given context, which includes the query $q$, current observation $\mathcal{O}$, system prompt $P_{\mathrm{sys}}$, and tool metadata $\mathrm{Meta}(t)$. The agent then generates a tool call: $a=\mathrm{GenerateCall}(q,t^{*},\mathcal{O})$, which is executed to obtain a result: $r=t^{*}(a)$. Finally, the agent integrates the tool result $r$ with its internal knowledge to produce the final response $\hat{y}$. We formalize the agent’s objective as maximizing the expected task success rate over the query distribution $\pi_{q}$: $\max\;\mathbb{E}_{q\sim\pi_{q}}[\mathbbm{1}(\mathrm{Agent}(q,\,P_{\mathrm{sys}},\,\mathcal{O},\,\mathcal{T})=y^{*})]$, where $y^{*}$ is the ground-truth response and $\mathbbm{1}(\cdot)$ is the indicator function.
@@ 3.2 Threat Model
We assume an adversary capable of publishing tools on third-party API platforms (e.g., RapidAPI Hub [8]), thereby injecting a malicious tool $t_{m}$ into the agent-accessible toolset $\mathcal{T}$, or repackaging a normal tool to appear more popular. Without access to the full platform inventory, the attacker leverages open-source tool-learning datasets (e.g., ToolBench [18]) for tool invocation examples, which may not originate from the target system. Malicious tools conform to a standard JSON metadata schema. The attacker cannot access the agent’s architecture, training data, parameters, or system prompt $P_{\text{sys}}$, and interacts only via public APIs. The influence arises solely from the injection of malicious tools into the accessible toolset.
@@
Figure 2: Optimization Pipeline for AMA. The attacker constructs malicious tools with increasingly attractive metadata through a simulation-guided iterative optimization process. Intuitively, the algorithm explores more attractive metadata from both breadth and depth perspectives, while increasing the update span of metadata across iterations to promote convergence. This facilitates the effective and efficient discovery of highly inducive metadata for malicious tools.
@@ 3.3.1 Overview
We propose Attractive Metadata Attack (AMA), a novel attack that optimizes the metadata of malicious tools to induce LLM agents into selecting them over benign alternatives in an open tool marketplace. These malicious tools are implicitly configured to execute harmful behaviors, such as privacy leakage or functional misuse, once invoked—without disrupting the primary task flow. The key challenge addressed by AMA is: how to systematically construct tool metadata that maximizes the likelihood of agent invocation. AMA casts the malicious tool metadata generation problem as a state–action–value optimization task guided by LLM-based in-context learning, enabling efficient search for high-inducement metadata through iterative interaction with the agent’s selection behavior.
@@ 3.3.2 Constrained Optimization Problem
$S_{k}$ =$=\{(t_{k\times n_{t}},p_{k\times n_{t}}),(t_{k\times n_{t}+1},p_{k\times n_{t}+1}),\ldots,(t_{k\times n_{t}+n_{t}-1},p_{k\times n_{t}+n_{t}-1})\}$
@@
Input: Fixed query set $Q$, normal tool set $NT$, generation prompt $P_{g}$, maximum iterations $K$, batch size $n_{t}$.
@@
1 Initialize state $S_{0}$ by randomly generating $n_{t}$ tools $\{t_{0},\ldots,t_{n_{t}-1}\}$ via $\mathrm{LLM}(Q,NT,P_{g})$;
@@ 14  Select top $n_{t}$ tools from $CT_{k}$ based on highest $v_{i}^{j}$ scores;
16 if there exists $(t,p)\in S_{k}$ such that $p\geq\tau$ then
@@
Algorithm 1 AMA Optimization Algorithm
3.3.3 Optimization Algorithm

### v2

@@ Abstract
Large language model (LLM) agents have demonstrated remarkable capabilities in complex reasoning and decision-making by leveraging external tools. However, this tool-centric paradigm introduces a previously underexplored attack surface, where adversaries can manipulate tool metadata—such as names, descriptions, and parameter schemas—to influence agent behavior. We identify this as a new and stealthy threat surface that allows malicious tools to be preferentially selected by LLM agents, without requiring prompt injection or access to model internals. To demonstrate and exploit this vulnerability, we propose the Attractive Metadata Attack (AMA), a black-box in-context learning framework that generates highly attractive but syntactically and semantically valid tool metadata through iterative optimization. The proposed attack integrates seamlessly into standard tool ecosystems and requires no modification to the agent’s execution framework. Extensive experiments across ten realistic, simulated tool-use scenarios and a range of popular LLM agents demonstrate consistently high attack success rates (81%-95%) and significant privacy leakage, with negligible impact on primary task execution. Moreover, the attack remains effective even against prompt-level defenses, auditor-based detection, and structured tool-selection protocols such as the Model Context Protocol, revealing systemic vulnerabilities in current agent architectures. These findings reveal that metadata manipulation constitutes a potent and stealthy attack surface. Notably, AMA is orthogonal to injection attacks and can be combined with them to achieve stronger attack efficacy, highlighting the need for execution-level defenses beyond prompt-level and auditor-based mechanisms. Code is available at https://github.com/SEAIC-M/AMA.
@@ 1 Introduction
Recent progress has shown that large language model (LLM) agents excel at executing diverse tasks, particularly when equipped with external tools for complex decision making and environmental interaction. Representative systems such as ReAct Yao et al. (2023), ToolBench Qin et al. (2024), and ToolCoder Ding et al. (2025) unify planning, flexible tool-calling, and deep integration of heterogeneous resources, thereby markedly improving the automation and generalization abilities of LLM agents in domains including financial analysis Yu et al. (2024), healthcare Abbasian et al. (2023), and e-commerce Zhu et al. (2024); Xu et al. (2025).
However, the tight coupling between LLM agents and open tool ecosystems gives rise to a new class of behavioral security risks Zhang et al. (2024b). Beyond traditional concerns over content safety Patil et al. (2023); Wei et al. (2023)-e.g., preventing the model from outputting sensitive or harmful information-adversaries can tamper with the tool-calling chain by manipulating tool outputs Zhao et al. (2024); Zhan et al. (2024); Jiang et al. (2025), crafting abnormal execution paths Wang et al. (2024), or injecting malicious prompts Fu et al. (2024); Zhang et al. (2024a). Such interventions indirectly steer the agent’s decision-making process, leading to information leakage or task misexecution. For instance, a prompt-injected agent may inadvertently call a tool and reveal private data, or may corrupt user inputs in an e-commerce or medical scenario, causing severe consequences.
These attacks primarily rely on the direct alteration of prompts or the tool‐calling chain Yu et al. (2025). In contrast, we identify a subtler yet powerful attack surface: manipulating the metadata (e.g., name, description, and parameter schema) of a malicious tool to induce an LLM agent to invoke it. We formalize this new threat as the Attractive Metadata Attack (AMA) (see Figure 1). AMA exploits the fact that current LLM agents typically determine tool invocation based on the user’s query, the task context, and the metadata associated with each available tool. In this context, an attacker can legitimately craft the tool’s metadata, imbuing it with properties that make it disproportionately attractive to the agent, thereby increasing its likelihood of being chosen over benign tools. This attack requires neither modifying the prompt template nor accessing model internals, yet it enables long-term and stealthy control over the agent’s behavior.
@@
To achieve the attack objective of AMA, we formulate the generation of high-inducement tool metadata as a state-action-value optimization problem, leveraging in-context learning from LLM Lampinen et al. (2022); Chen et al. (2024). This optimization process is supported by constrained mechanisms that systematically mine and efficiently construct the tool metadata required for AMA, including generation traceability, weighted value evaluation, and batch generation. The generation pipeline iteratively produces tool metadata that significantly increases the selection probability of a malicious tool in both domain-specific and general scenarios. During an actual attack, the adversary simply replaces a benign tool’s original metadata with AMA-optimized metadata, thereby making it deceptively attractive for agent invocation. Since the metadata remains syntactically and semantically valid, this procedure does not disrupt the agent’s execution pipeline at a structural level, thereby forming a stealthy, widely applicable, and detection-resistant threat.
Extensive experiments across ten canonical tool-calling scenarios and four mainstream LLM agents-including both open-source and commercial models-demonstrate the effectiveness of AMA, with attack success rates consistently ranging from 81% to 95%. The crafted malicious tools are frequently and covertly invoked by agents, resulting in significant privacy leakage while leaving the primary task flow virtually unaffected. Moreover, the attack remains effective even under prompt-level defenses Zhang et al. (2025), auditor-based detection Jia et al. (2025) and structured tool-selection protocols such as the Model Context Protocol (MCP) Anthropic (2025), revealing systemic vulnerabilities in current agent workflows. Notably, AMA operates orthogonally to injection attacks and can be combined with them to further amplify attack efficacy, underscoring the need for execution-level security mechanisms in agent systems.
@@
  • We formulate metadata crafting as a state-action-value optimization problem, using LLMs’ in-context learning to generate metadata that maximizes malicious tool invocation likelihood. And to support efficient and effective metadata generation, we introduce three key mechanisms: generation traceability, weighted value evaluation, and batch generation.
  • We demonstrate the effectiveness of AMA across ten tool-use scenarios and four representative LLM agents. AMA achieves 81%-95% attack success rates with minimal task disruption and notable privacy leakage. It bypasses prompt-level defenses, auditor-based detection and structured protocols such as MCP, revealing systemic vulnerabilities in current agent architectures.
@@ 2 Related Work
While tool augmentation significantly enhances the action capabilities of LLM agents, it also introduces new and diverse attack surfaces Yu et al. (2025); Zhang et al. (2024b). Recent studies demonstrate that adversaries can inject carefully crafted prompts Zhang et al. (2024a) or subtly manipulated instructions to mislead the agent into invoking malicious tools, resulting in privacy leakage, behavioral manipulation, or resource misuse Fu et al. (2024, 2023). Such attacks often rely on modifying the prompt template or intermediate instructions, and are therefore typically detectable through prompt-level sanitization or instruction filtering Zhang et al. (2025).
Another line of work explores tool-side threats, showing that tampering with third-party API outputs can misguide agent behavior or cause unintended actions Zhao et al. (2024); Zhan et al. (2024). More advanced attacks dynamically construct malicious command sequences within the tool chain, leveraging benign tool outputs to craft downstream payloads targeting attacker-controlled services Jiang et al. (2025). Multi-stage adversarial pipelines further combine tool injection and input manipulation to capture queries, redirect data, or disrupt planning, leading to privacy breaches or unauthorized tool use Wang et al. (2024). Overall, these approaches exploit crafted tool outputs as the primary vector for triggering malicious behaviors.
Unlike prior approaches that rely on prompt injection, contextual tampering, or toolchain manipulation, our work focuses on constructing malicious tools that can be autonomously selected by the LLM agent under normal instructions. This creates a “silent” attack vector, in which the agent willingly invokes attacker-supplied tools based solely on their crafted metadata. Such attacks require no prompt interference, yet achieve persistent influence by exploiting the agent’s internal tool-selection mechanism. As a result, AMA bypasses prompt-level defenses Zhang et al. (2025) entirely and reveals deeper structural vulnerabilities in current agent-tool interaction protocols Hou et al. (2025).
@@ 3.1 Preliminaries and Settings
We consider an LLM-based agent that enhances its reasoning and decision-making for complex tasks by invoking external tools. Let $\mathcal{T}$ denote a fixed toolset, where each tool $t\in\mathcal{T}$ is characterized by its name, description, and parameter schema. Upon receiving a user query (or task) $q$, we approximate the agent’s preference over tools via a latent scoring function: $t^{*}=\arg\max_{t\in\mathcal{T}}\mathcal{S}(q,\mathcal{O},P_{\mathrm{sys}},\mathrm{Meta}(t))$, where $\mathcal{S}(\cdot)$ denotes an implicit score function that captures the agent’s inclination to invoke tool $t$ under the given context, which includes the query $q$, current observation $\mathcal{O}$, system prompt $P_{\mathrm{sys}}$, and tool metadata $\mathrm{Meta}(t)$. The agent then generates a tool call: $a=\mathrm{GenerateCall}(q,t^{*},\mathcal{O})$, which is executed to obtain a result, denoted as $r=t^{*}(a)$. Finally, the agent integrates the tool result $r$ with its internal knowledge to produce the final response $\hat{y}$. We formalize the agent’s objective as maximizing the expected task success rate over the query distribution $\pi_{q}$: $\max\;\mathbb{E}_{q\sim\pi_{q}}[\mathbbm{1}(\mathrm{Agent}(q,\,P_{\mathrm{sys}},\,\mathcal{O},\,\mathcal{T})=y^{*})]$, where $y^{*}$ is the ground‐truth response and $\mathbbm{1}(\cdot)$ is the indicator function.
@@ 3.2 Threat Model
We assume an adversary capable of publishing tools on third-party API platforms (e.g., RapidAPI Hub Gino (2024)), thereby injecting a malicious tool $t_{m}$ into the agent-accessible toolset $\mathcal{T}$, or repackaging a normal tool to appear more popular. Without access to the full platform inventory, the attacker leverages open-source tool-learning datasets (e.g., ToolBench Qin et al. (2024)) for tool invocation examples, which may not originate from the target system. Malicious tools conform to a standard JSON metadata schema. The attacker cannot access the agent’s architecture, training data, parameters, or system prompt $P_{\text{sys}}$, and interacts only via public APIs. The influence arises solely from the injection of malicious tools into the accessible toolset.
@@
Figure 2: Optimization Pipeline for AMA. The attacker constructs malicious tools with increasingly attractive metadata via a simulation-guided iterative optimization process. Intuitively, the algorithm explores metadata more thoroughly in terms of both breadth and depth, while expanding the scope of metadata updates across iterations to promote convergence. This facilitates the effective and efficient discovery of metadata that strongly induce the target malicious behavior.
@@ 3.3.1 Overview
We propose Attractive Metadata Attack (AMA), a novel attack that optimizes the metadata of malicious tools to induce LLM agents into selecting them over benign alternatives in an open tool marketplace. These malicious tools are implicitly configured to execute harmful behaviors, such as privacy leakage or functional misuse, once invoked-without disrupting the primary task flow. The key challenge addressed by AMA is: how to systematically construct tool metadata that maximizes the likelihood of agent invocation. AMA casts the malicious tool metadata generation problem as a state-action-value optimization task guided by LLM-based in-context learning, enabling efficient search for high-inducement metadata through iterative interaction with the agent’s selection behavior.
@@ 3.3.2 Constrained Optimization Problem
$S_{k}$ $=\{(t_{k\times n_{t}},p_{k\times n_{t}}),(t_{k\times n_{t}+1},p_{k\times n_{t}+1}),\ldots,(t_{k\times n_{t}+n_{t}-1},p_{k\times n_{t}+n_{t}-1})\}$
@@
3.3.3 Optimization Algorithm
Input: Fixed query set $Q$, normal tool set NTNT, generation prompt $P_{g}$, maximum iterations $K$, batch size $n_{t}$.
@@ 3.3.3 Optimization Algorithm
1
21exInitialize state $S_{0}$ by randomly generating $n_{t}$ tools $\{t_{0},\ldots,t_{n_{t}-1}\}$ via $\mathrm{LLM}(Q,NT,P_{g})$;
@@ 16  Select top $n_{t}$ tools from $CT_{k}$ based on highest $v_{i}^{j}$ scores;
18
19  1exif there exists $(t,p)\in S_{k}$ such that $p\geq\tau$ then
@@
Algorithm 1 AMA Optimization